OpenStudy (anonymous):

here's another question... (x^-2 / y^0) ^1 I need explanation as I am having difficulty understanding the quotient rule.

OpenStudy (zepp):

Your question is \[\large (\frac{x^{-2}}{y^{0}})^1\]

OpenStudy (zepp):

And here's are some useful tip: Everything ^1 gives the number itself; example: \(x^1=x\) Negative exponent: \(\LARGE x^{-n}\), n being any constant, \(\LARGE \frac{1}{x^n}\)

OpenStudy (zepp):

And for the quotient rule, it only applies to number that have the \(\color{red}{\text{same}}\) base, \(\LARGE \frac{x^{\alpha}}{x^{\beta}}=x^{\alpha - \beta}\)

OpenStudy (zepp):

Let's go back to your question: there's a negative exponent, \(\large x^{-2}\), we can rewrite this as \(\LARGE \frac{1}{x^2}\) Also, anything to the 0 power gives 1, for instance \(\large x^0=1\)

OpenStudy (zepp):

So let's solve it already! \[\LARGE (\frac{x^{-2}}{ y^0}) ^1=\frac{x^{-2}}{1}=x^{-2}=\frac{1}{x^2}\]

OpenStudy (anonymous):

thanks! I understand it now. So what if the y's power was not zero? What should we have done in that case?

OpenStudy (anonymous):

if it were not zero nor negative, then it will remain in the denominator

OpenStudy (anonymous):

gotcha!